OpenStudy (anonymous):

Solve the proportion. 2/15 = x/60 a. 450 b.8 c.4 d.30

OpenStudy (coconutjj):

use cross multiplication

OpenStudy (anonymous):

cross your heart \[\frac{2}{15}=\frac{x}{60}\\ x=\frac{60\times 2}{15}\] cancel first, multiply last

OpenStudy (anonymous):

I got 8

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

same :D